OpenStudy (anonymous):

4x+8/2x + X-1/3x = PLEASE HELP!! PLEASE.

OpenStudy (aum):

In 8/2x is x in the denominator?

OpenStudy (anonymous):

I know this is simple but for some reason im stuck. :/

OpenStudy (anonymous):

no its 4x+8 -------over 2x

OpenStudy (aum):

You should use parenthesis to group things together. Is it (4x+8) / (2x) + (x-1) / (3x) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry ill do that next time. Its smart..

OpenStudy (aum):

\[\frac{4x+8}{2x} + \frac{x-1}{3x} = \frac{2(2x+4)}{2x} + \frac{x-1}{3x} = \frac{2x+4}{x} + \frac{x-1}{3x} \]

OpenStudy (aum):

The least common denominator is 3x. \[\frac{2x+4}{x} + \frac{x-1}{3x} = \frac{6x+12}{3x} + \frac{x-1}{3x} = \frac{7x+11}{3x}\]

OpenStudy (anonymous):

So then the final anser would be (7x+11)/(3x). Thats the simplest form, right?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Okay thank you so much. You really helped me out. :)

OpenStudy (aum):

You are welcome.